three hundred tickets were sold at a school play. students paid 150 each and guests paid 350 each. If the total receipts for the play were 75,000, how many tickets of each kind were sold?
Answer by mananth(16946)

x= number of student tickets
y= number of guest tickets

1 x + 1 y = 300 .............1
Total value
150 x + 350 y = 75000 .............2
Eliminate y
multiply (1)by -350
Multiply (2) by 1
-350 x -350 y = -105000
150 x + 350 y = 75000
Add the two equations
-200 x = -30000
/ -200
x = 150
plug value of x in (1)
1 x + 1 y = 300
150 + y = 300
y = 300 -150
y = 150
y = 150
x= 150 number of student tickets
y= 150 number of guest tickets
m.ananth@hotmail.ca